OpenStudy (anonymous):

what is the domain and range of this f(x)=1/3x-2

OpenStudy (anonymous):

x=-1/2

OpenStudy (anonymous):

....

OpenStudy (anonymous):

x cannot = 2/3

OpenStudy (anonymous):

{XER: x cannot = 2/3}

OpenStudy (anonymous):

If the function is (1/3x) - 2 then the domain is all real numbers other than 0 and the range is all real numbers other than -2. If the function is 1/(3x-2) then the domain is all real numbers other than 2/3 and the range is all real numbers other than 0